OpenStudy (anonymous):

evaluate the indefinite integral of dx/8x^3

OpenStudy (anonymous):

\[\bf \int\limits_{}^{}\frac{ dx }{ 8x^3 }=\frac{ 1 }{ 8 }\int\limits_{}^{}\frac{ dx }{ x^3 }=\frac{ 1 }{ 8 }\int\limits_{}^{}x^{-3}dx=?\]Hint:\[\bf \int\limits_{}^{}x^n dx=\frac{ x^{n+1 }}{ n+1 }+C\]@91004775

OpenStudy (anonymous):

@91004775

OpenStudy (anonymous):

i got 1/3 (x^-2/-2)